OpenStudy (anonymous):

Help with Algebra 2, please! Fanning and giving medal. (: I've been stuck on this question for about a week now (It's about linear programming which I'm a total noob at), and it's about candles. (Brick and Egyptian candles) Question along with helpful hints are in the comments. Any help is appreciated. ^_^

OpenStudy (anonymous):

3. You ask the production team to make some candles for the sale. The team is not sure how many they can produce in your short time frame, so they ask for an acceptable range for how many of each candle they should make. Use your costs and recommended retail prices to write a linear programming model to show the executives possible profits for selling certain numbers of each type of candle during this promotion. Decide on the constraints for the numbers of candles made for the sale and explain your reasoning for these constraints. Graph the feasible region and find maximum and minimum profits. COSTS & RETAIL PRICES: I have a price margin of about $5. The price for a brick candle would be about $6.10 (wax cost is about 0.0075, so 600 * 0.0075 = 4.5 + 1.6 = 6.1), while the price for the Egyptian candle would be $4.30 (0.0075 * 360 = 2.7 + 1.6 = 4.3). If I was a customer and wanted the brick candle, I’d pay $11.10, or if I wanted the Egyptian candle, I’d pay $9.30.

OpenStudy (dumbcow):

what are the time constraints?? the profit function is 5 times total candles --> P = 5E + 5B

OpenStudy (anonymous):

Not sure what numbers I should choose for the constraints tbh. .-. @dumbcow